the position of a particle is given by
r = ( t^2 – 4) i + (t – 4) j m
Find the equation of the trajectory of the particle

Grade:11

1 Answers

Vikas TU
7 years ago
r = xi + yj
comparing with the given r vector
x = t^2 + 4....................(1)
and y = t – 4 
or t = y + 4.... put it in eqn. (1)
x  = (y+4)^2 + 4
is the required trajectory eqn.